# Mapple: A Domain-Specific Language for Mapping Distributed Heterogeneous Parallel Programs

### v1

@@ 6.2. Performance Tuning
We demonstrate that Mapple facilitates effective performance tuning by enabling the creation of mappers that outperform expert-designed C++ mappers. As shown in Table 2, Mapple achieves up to 1.34$\times$ speedup over the baseline C++ implementations. Even with a simple search-based algorithm, Mapple uncovers substantial optimization opportunities, helping users discover mappers that achieve higher throughput.
We analyze the sources of the observed performance improvements. For the six matrix multiplication applications (indexed 4 to 9), the speedup is entirely due to tuning the index mappings, although other features of Mapple contribute to code size reduction. In contrast, for the three scientific applications, the performance gains stem from assigning memories differently than the expert-written mappers—a capability enabled by Mapple’s full feature set, as discussed in Section 7.1.
@@
Mapple Tuned Speedup 1.34$\times$ 1.02$\times$ 1.04$\times$ 1.09$\times$ 1.09$\times$ 1.09$\times$ 1.09$\times$ 1.07$\times$ 1.31$\times$
We also present a case study highlighting the critical role of precise control over mapping strategies. Our investigation demonstrates significant performance variations resulting from minor alterations in the hierarchical_block2D function for the Cannon’s, PUMMA, and SUMMA algorithms. Specifically, we contrast the mapping function specified by the algorithm with an alternative approach integrating runtime heuristics: the runtime system dynamically assigns the iteration point to one of the four GPUs on the node, selecting the GPU with the least workload at runtime, rather than adhering to a predetermined distribution.
The throughput result depicted in Figure 13 underscores the pronounced performance discrepancies between the two approaches. The x-axis indicates different machine sizes, and the y-axis indicates the throughput per node. The “Algorithm Specification” line shows the throughput achieved by correctly implementing the corresponding algorithms, which match the performance results reported in prior work. The “Runtime Heuristics” line shows the throughput achieved by the alternative approach. In the 1-node scenario, the slowdown can reach up to 3.5$\times$. This disparity stems from the additional data movement induced by divergent mapping decisions. Furthermore, the runtime heuristics-based mapping leads to out-of-memory (OOM) errors on 32-GPU runs for the PUMMA and SUMMA algorithms, because mapping decisions influence where the data is physically materialized in memory. This example shows that it is important to precisely control the mapping to minimize data movement and optimize memory resource utilization.
(a) Cannon’s
(b) PUMMA
(c) SUMMA
Figure 13. Throughput comparison between the original hierarchical_block2D mapping function (specified by the Cannon’s, PUMMA, and SUMMA algorithm) and a runtime heuristics-based mapper. The runtime heuristics-based mapping can lead to out-of-memory error (denoted by “OOM”).
@@
Figure 15. Geometric mean of improvement percentage w.r.t. aspect ratios of iteration space.
Figure 16. Geometric mean of improvement percentage w.r.t. area of iteration space per node.
@@ 6.3. Performance Study of the Decompose Primitive
To investigate the impact of the decompose primitive, we compare it against the default heuristics for reshaping machine spaces from Algorithm 1. We evaluate the end-to-end performance of both mappers on stencil-pattern applications, which are a dominant communication pattern in image processing and scientific computing. Each point in the iteration space updates its value based on its nearest neighbors, making these applications highly sensitive to how computations are partitioned and mapped. As a result, demonstrating effectiveness on stencil applications provides strong evidence of the practical benefits of decompose for reducing communication in widely used computational workloads.
Table 3. Parameter space in performance comparison
Parameter Values
Aspect ratio of total iteration space ($x:y$) $1:1,1:2,1:4,1:8,1:16,1:32$
Area of iteration space per node ($x*y\ /\ \#nodes)$ $10^{6},10^{7},10^{8},2\times 10^{8},4\times 10^{8}$
Number of GPUs 4, 8, 16, 32, 64, 128
To ensure a fair comparison, we systematically vary the 2D stencil parameters as shown in Table 3. We test six aspect ratios from $1:1$ to $1:32$, reflecting shapes commonly seen in scientific simulations (e.g., liquid films (Nave et al., 2010), shock tubes (Wong et al., 2019)). As discussed in Section 4.1, Algorithm 1 can lead to suboptimal results. We also vary the iteration space per node across five sizes ($10^{6}$ to $4\times 10^{8}$) to capture different communication-to-computation ratios, and scale the number of GPUs from 4 to 128 in powers of two. In total, we evaluate 180 configurations.
We report the percentage improvement, with the distribution shown in Figure 17, ranging from 0% to 83%. The geometric mean improvement is 16%. We also plot the geometric mean improvement across varying values of each parameter. In Figure 17, the y-axis shows the geometric mean improvement percentage, and the x-axis shows the iteration space aspect ratio. As the ratio increases from $1:1$ to $1:32$, the improvement rises from 7% to 27%, highlighting the growing advantage of mappers using the decompose primitive. This trend is expected, as Algorithm 1 evenly splits processors—a strategy that performs well when the iteration space is nearly square.
Figure 17 shows how the improvement percentage varies with the iteration space size per node. As the area increases from $10^{6}$ to $4\times 10^{8}$, the improvement drops from 32% to 5%. Larger iteration spaces reduce the communication-to-computation ratio, diminishing the relative benefit of minimizing communication. Figure 17 plots improvement against the number of nodes and GPUs. The improvement peaks at 26% on 4 nodes (16 GPUs). From 1 to 4 nodes, growing inter-node communication enhances the benefit of reducing bandwidth-bound traffic. Beyond 4 nodes, inter-rack latency becomes the bottleneck, diminishing the impact of improved mappings.
@@ 7. Discussion
7.1. Additional Performance Optimization Features in Mapple
@@
[…]
$\quad\ \ \ \text{IndexTaskMap TaskName var}\mid\text{...}$
[…]
Figure 18. Mapple’s Grammar
@@
7.2. Generalization of the Decompose Primitive
While Section 4 addresses isotropic communication across iteration space dimensions, the decompose primitive can be extended to anisotropic patterns, such as uneven halo widths and dimension-specific all-to-all exchanges (e.g., for data transposes). This generalization supports a much wider range of computations on block-structured grids. In both cases, only the objective function changes; the same minimization procedure from Section 4.3 still applies.
7.2.1. Anisotropic halo communication
To generalize, we extend the communication area from Section 4.2 to a volume that captures the total data exchanged between distributed machines. In isotropic cases, where halo width is uniform, volume reduces to area. For anisotropic patterns, varying halo width across dimensions must be considered. Letting $h_{n}$ denote the width in dimension $n$, the total communication volume $V$ in a $k$-dimensional space is $V=\sum_{n=1}^{k}d_{i_{n}}h_{n}\prod_{m\neq n}l_{m}$. For $l_{n}=w_{n}d_{i_{n}}$, the volume, $V$, can also be expressed as $V=\left(\sum_{n=1}^{k}\frac{h_{n}}{w_{n}}\right)\left(\prod_{m=1}^{k}l_{m}\right)$, where $\prod_{m=1}^{k}l_{m}$ is constant for a certain iteration space.
7.2.2. Transpose via all-to-all communication along certain dimensions
@@ 8.1. Task-based Programming Systems
Legion (Bauer et al., 2012)
StarPU (Augonnet et al., 2010)
Chapel (Chamberlain et al., 2007)
Charm++ (Kale and Krishnan, 1993)
PaRSEC (Danalis et al., 2015)
X10 (Charles et al., 2005)
HPX (Heller et al., 2017)
OpenMP (Chandra et al., 2001)
OmpSs (Duran et al., 2011)
@@
Table 4 summarizes mapping features in HPC systems, all of which are supported by Mapple. Systems like Legion (Bauer et al., 2012) and StarPU (Augonnet et al., 2009) offer low-level C/C++ mapping interfaces via asynchronous callbacks, requiring deep familiarity with internal runtime abstractions. Other HPC systems offer less direct mapping control. ZPL (Deitz et al., 2004), Chapel (Chamberlain et al., 2007), and X10 (Charles et al., 2005) expose predefined data distributions (e.g., blocked, cyclic). HPX (Heller et al., 2017) and PaRSEC (Danalis et al., 2015) rely on schedulers with tunable policies for heterogeneous execution. OpenMP (Chandra et al., 2001), OmpSs (Duran et al., 2011), Charm++ (Kale and Krishnan, 1993), and Chapel’s low-level APIs allow programmers to specify task placement directly in the source code, ranging from processor annotations (OpenMP, OmpSs) to custom C++ mapping classes (Charm++). Task-based systems in data analytics and AI such as MapReduce (Dean and Ghemawat, 2008), Spark (Zaharia et al., 2010), Dask (Rocklin, 2015), Ray (Moritz et al., 2018), and Pathways (Barham et al., 2022) rely on runtime mapping heuristics. This eases development but limits performance tuning when heuristics fall short. Mapple offers similar ease of use while enabling fine-grained control.
@@ 8.2. Loop Transformations
A large body of work on compiler algorithms uses loop transformations to maximize parallelism and minimize communication, including unimodular transformations (Banerjee et al., 1990; Banerjee, 2007; Wolf, 1992; Wolf and Lam, 1991b; Wolfe, 1982) (i.e., combinations of loop interchange, reversal, and skewing), loop block/tiling, fusion, fission, reindexing, and scaling (Lim et al., 1999, 2001; Wolf and Lam, 1991a; Xue, 2000), and affine transformations used by the polyhedral model (Bondhugula, 2013; Bondhugula et al., 2008; Acharya et al., 2018). These works share a similar motivation to Mapple, though the goals and specific transformations are different as explained in Section 1.
Another line of work uses scheduling languages (Chen et al., 2008; Ragan-Kelley et al., 2012; Zhang et al., 2018; Baghdadi et al., 2019; Kjolstad et al., 2017; Senanayake et al., 2020; Chen et al., 2018a; Yadav et al., 2022b) to separate the algorithm from the schedule, based on traditional loop optimizations such as split, collapse, and loop reordering. This separation enables programmers to just change the schedule when moving to different hardware. A growing body of research focuses on auto-scheduling (Adams et al., 2019; Chen et al., 2018b; Mullapudi et al., 2016; Zheng et al., 2020, 2022; Tollenaere et al., 2023) to reduce the manual effort required in writing schedules. Task-based systems share a similar approach in separating the algorithm from its schedule or mapping. However, prior work on scheduling languages does not address the challenges of mapping iteration spaces onto processor spaces of differing dimensionality in distributed settings.
@@ 9. Conclusion
We present Mapple, a high-level interface that simplifies the development of mappers for task-based parallel systems. Mapple overcomes key limitations of existing mapping interfaces by abstracting low-level runtime details while exposing transformation primitives to resolve dimensionality mismatches between iteration and processor spaces. At the core of Mapple is the decompose primitive, which minimizes communication volume based on theoretical analysis. Across nine distributed applications, Mapple reduces mapper code size by 14$\times$ and achieves up to 1.34$\times$ speedup over expert-written C++ mappers. Moreover, the decompose primitive outperforms existing heuristics by up to 1.83$\times$. These results demonstrate the practical effectiveness of Mapple in enabling the development of concise, high-performance mappers for distributed applications.

### v2

@@ 6.2. Performance Tuning
[…]
This performance gain comes from Mapple’s high-level declarative design, which is still expressive for high-performance mapping. Although our compiler translates Mapple mappers into standard C++ code, making the same strategies theoretically possible in C++, experts have not discovered them in practice. This is likely due to the low-level complexity of C++ mappers, where exploring new strategies often requires rewriting hundreds of lines of code. In contrast, similar changes in Mapple take only a few lines, making tuning much easier.
We further analyze the sources of performance gains. For the six matrix multiplication applications (indexed 4 to 9), improvements arise solely from optimized index mappings. For the three scientific applications, they result from different memory assignments enabled by Mapple’s full feature set (Section 7). Finally, case studies in Section A.5 (see supplementary material) illustrate how mapping choices impact performance.
@@
[…]
@@
Figure 13. Geometric mean of improvement percentage w.r.t. aspect ratios of task spaces.
Figure 14. Geometric mean of improvement percentage w.r.t. area of task space per node.
@@ 6.3. Performance Study of the Decompose Primitive
To evaluate the impact of the decompose primitive, we compare it with the default machine space reshaping heuristic from Algorithm 1 (see supplementary material). We measure end-to-end performance on stencil applications, which are common in image processing and scientific computing, where each point updates based on its neighbors. As these applications are sensitive to data partitioning and mapping, the results provide strong evidence of decompose improving communication efficiency in practical workloads.
To ensure a fair comparison, we systematically vary the 2D stencil parameters. We test six aspect ratios ($1:1$, $1:2$, $1:4$, $1:8$, $1:16$, $1:32$), representative of shapes typical in scientific simulations (e.g., liquid films (nave2010direct), shock tubes (wong2019high)). We also vary the task space size per node across five scales ($10^{6}$, $10^{7}$, $10^{8}$, $2{\times}10^{8}$, $4{\times}10^{8}$) for different communication-to-computation ratios, and scale the number of GPUs from 4 to 128 in powers of two. In total, we evaluate 180 configurations.
We report percentage improvements in Figure 15, ranging from 0% to 83%, with a geometric mean of 16%. Figure 15 plots the geometric mean improvement against the task space aspect ratio. As the ratio increases from $1:1$ to $1:32$, the improvement rises from 7% to 27%, showing the growing benefit of using the decompose primitive. This trend is expected since Algorithm 1 evenly splits processors, which works best when the task space is nearly square.
Figure 15 shows how improvement varies with task space size per node. As the area grows from $10^{6}$ to $4\times 10^{8}$, the improvement decreases from 32% to 5%, since larger spaces lower the communication-to-computation ratio. Figure 15 plots improvement against machine size. The benefit peaks at 26% on 4 nodes: increasing from 1 to 4 nodes amplifies inter-node communication, while beyond 4 nodes, inter-rack latency becomes the bottleneck, reducing the gain.
@@ 7. Discussion
Program $::=\;\text{Statement}^{+}$
Statement $::=\;\text{TaskMap}\mid\text{DataMap}\mid\text{DataLayout}\mid{}$
$\hskip 17.00024pt\text{FuncDef}\mid\mathrm{IndexTaskMap}\;\mathit{t}\;\mathit{v}\mid\cdots$
TaskMap $::=\;\mathrm{Task}\;\mathit{t}\;\text{Proc}^{+}$
DataMap $::=\;\mathrm{Region}\;\mathit{t}\;\mathit{r}\;\text{Proc}\;\text{Memory}^{+}$
Proc $::=\;\mathrm{CPU}\mid\mathrm{GPU}\mid\mathrm{OMP}$
Memory $::=\;\mathrm{SYSMEM}\mid\mathrm{FBMEM}\mid\mathrm{ZCMEM}$
DataLayout $::=\;\mathrm{Layout}\;\mathit{t}\;\mathit{r}\;\text{Proc}\;\text{Constraint}^{+}$
Constraint $::=\;\mathrm{SOA}\mid\mathrm{AOS}\mid\mathrm{C\_order}\mid\mathrm{F\_order}\mid\mathrm{Align}{\ ==\ }\mathit{i}$
FuncDef $::=\;\mathrm{def}\;\mathit{v}(\mathit{v}^{+}):\;\text{FuncStmt}^{+}$
FuncStmt $::=\;\mathit{v}\;{=}\;\text{Expr}\mid\mathrm{return}\;\text{Expr}$
Expr $::=\;\mathit{v}\mid\mathit{v}(\text{Expr}^{+})\mid\mathrm{Machine}(\text{Proc})\mid{}$
$\hskip 17.00024pt\text{Expr}\mathbin{.}\text{Expr}\mid\text{Expr}\;\mathrm{Op}\;\text{Expr}\mid(\text{Expr})\mid{}$
@@
Primitive
Primitive $::=\;\mathrm{split}\mid\mathrm{merge}\mid\mathrm{swap}\mid\mathrm{slice}\mid\mathrm{decompose}$
$\mathit{t}\in\text{{Task Names}}\quad\mathit{r}\in\text{{Region Names}}\quad\mathit{v}\in\text{{Variables}}\quad\mathit{i}\in\text{{Integers}}$
Figure 16. Syntax of the Mapple language.
@@
While Mapple captures a broad range of optimization dimensions, it is intentionally constrained to focus on performance-critical mapping decisions that, in our experience, are the most common in practice. One current limitation is that Mapple is stateless, making it less suitable for highly dynamic workloads or computations that would benefit from retaining historical information across mapping decisions. Although it could be extended with a persistent state, we have not yet found this necessary. Beyond this, we have not encountered cases where Mapple is less expressive than the C++ interface. Notably, every mapper that achieved higher performance than the expert-written C++ implementations was expressible in Mapple, demonstrating that Mapple is effective despite being more restrictive than general-purpose C++, at least for the applications we studied.
@@ 8.1. Task-based Programming Systems
Legion (bauer2012legion)
StarPU (augonnet2010starpu)
Chapel (chamberlain2007parallelchapel)
Charm++ (kale1993charm++)
PaRSEC (danalis2015parsec)
X10 (charles2005x10)
HPX (heller2017hpx)
OpenMP (chandra2001parallelopenmp)
OmpSs (duran2011ompss)
@@
Table 3 summarizes mapping features in HPC systems, all of which are supported by Mapple. Systems like Legion (bauer2012legion) and StarPU (augonnet2009starpu) offer low-level C/C++ mapping interfaces via asynchronous callbacks, requiring deep familiarity with internal runtime abstractions. Other HPC systems offer less direct mapping control. ZPL (deitz2004abstractionszpl), Chapel (chamberlain2007parallelchapel), and X10 (charles2005x10) expose predefined data distributions (e.g., blocked, cyclic). HPX (heller2017hpx) and PaRSEC (danalis2015parsec) rely on schedulers with tunable policies for heterogeneous execution. OpenMP (chandra2001parallelopenmp), OmpSs (duran2011ompss), Charm++ (kale1993charm++), and Chapel’s low-level APIs allow programmers to specify task placement directly in the source code, ranging from processor annotations (OpenMP, OmpSs) to custom C++ mapping classes (Charm++). Task-based systems in data analytics and AI such as MapReduce (dean2008mapreduce), Spark (zaharia2010spark), Dask (rocklin2015dask), Ray (moritz2018ray), and Pathways (barham2022pathways) rely on runtime mapping heuristics. This eases development but limits performance tuning when heuristics fall short. Mapple offers better ease-of-use than low-level mapping interfaces while enabling finer-grained control than systems that rely entirely on heuristics built into the runtime system.
@@ 8.2. Loop Transformations
A large body of work on compiler algorithms uses loop transformations to maximize parallelism and minimize communication, including unimodular transformations (banerjee1990unimodular; banerjee2007loop; wolf1992improving; wolf1991loop; wolfe1982optimizing) (i.e., combinations of loop interchange, reversal, and skewing), loop block/tiling, fusion, fission, reindexing, and scaling (lim1999affine; lim2001blocking; wolf1991data; xue2000loop), and affine transformations used by the polyhedral model (bondhugula2013compiling; bondhugula2008practical; acharya2018polyhedral). These works share a similar motivation to Mapple, though the goals and specific transformations are different as explained in Section 1.
Another line of work uses scheduling languages (chen2008framework; ragan2012decoupling; zhang2018graphit; baghdadi2019tiramisu; kjolstad2017tensor; senanayake2020sparse; chen2018tvm; yadav2022spdistal) to separate the algorithm from the schedule, based on traditional loop optimizations such as split, collapse, and loop reordering. This separation enables programmers to just change the schedule when moving to different hardware. A growing body of research focuses on auto-scheduling (adams2019learning; chen2018learning; mullapudi2016automatically; zheng2020ansor; zheng2022amos; tollenaere2023autotuning) to reduce the manual effort required in writing schedules. Task-based systems share a similar philosophy of separating the algorithm from its schedule or mapping. However, prior scheduling languages such as DISTAL (yadav2022spdistal) and Distributed Halide (denniston2016distributed) are specialized compiler frameworks for tensor algebra that lower high-level tensor expressions to distributed execution. In contrast, Mapple addresses the general mapping problem in task-based runtime systems (see Section 7), supporting a broader class of task-based programs. Its design centers on the decompose primitive (see Section 4), which automatically resolves dimensionality mismatches between task and processor spaces. In our evaluation, DISTAL’s C++ mappers serve as baselines; Mapple reduces their code size by over 11× and discovers mappers that achieve higher performance.
@@ 9. Conclusion
[…]
Appendix A Supplementary Material
A.1. Example of Cannon’s Algorithm
Cannon’s algorithm computes $A=B\cdot C$ by arranging all processors in a 2D grid, and assigning a tile of each matrix to each of the processors. Every processor $p$ alternates between systolic shifts of the tiles of $B$ and $C$ with $p$’s row and column neighbors and accumulating the product of $p$’s current $B$ and $C$ tiles into $A$.
We illustrate the task-based pseudocode for Cannon’s algorithm, adapted from code generated by DISTAL (yadav2022distal). The program in Figure 1(a) partitions the input matrices into tiles and launches the tiles task across the processor grid. Each tiles task, in turn, launches multiple systolic tasks that fetch tiles of $B$ and $C$ and accumulate their products into the corresponding tile of $A$. This program builds a task graph for distributed matrix multiplication, where data dependencies between tasks are automatically managed by the runtime system. The task space of the tiles task is two-dimensional, defined as $(i,j)\in\{0,\ldots,p_{x}-1\}\times\{0,\ldots,p_{y}-1\}$, where each point corresponds to a tile of the output matrix. The mapping from this task space to the processor space is specified by the mapper declared in the IndexTaskMap statement shown in Figure 1(b).
1task systolic(A: region(double),
2 B: region(double),
3 C: region(double))
4 writes(A), reads(A, B, C):
5 # Internally use cuBLAS for single-GPU matmul.
6 A += matmul(B, C)
7
8task tiles(A: region(double),
9 B: region(double),
10 C: region(double),
11 pB: partition(B),
12 pC: partition(C),
13 px: int, i: int, j: int)
14 writes(A), reads(A, B, C):
15 for k in (0, px):
16 k = (k + i + j) % px
17 systolic(A, pB[i, k’], pC[k’, j])
18
19task cannons_matmul(A: region(double),
20 B: region(double),
21 C: region(double),
22 px : int, py : int)
23 writes(A), reads(A, B, C):
24 pA, pB, pC = ... # Partition into px * py tiles.
25 for i, j in ({0, 0}, {px, py}):
26 tiles(pA[i, j], B, C, pB, pC, px, i, j)
(a) Application: pseudocode for Cannon’s algorithm.
1# Define a block mapping for a 2D launch domain.
2def block(Task task):
3 # ispace represents 2D launch domain of a x b.
4 is = task.ispace
5 # ipoint is an index point within launch domain.
6 ip = task.ipoint
7 # Get all GPUs as m nodes x n GPUs grid.
8 machine = Machine(GPU)
9 # Tile nodes; 3-dim grid: (m1 x m2) x n.
10 m3d = machine.decompose(0, is)
11 # Tile GPUs; 4-dim grid: (m1 x m2) x (n1 x n2).
12 m4d = m3d.decompose(2, is / m3d[:-1])
13 # Project the index point onto the 4D machine.
14 upper = (ip[0] * m4d[0] / is[0],
15 ip[1] * m4d[1] / is[1])
16 lower = (ip[0] % m4d[2], ip[1] % m4d[3])
17 # Unpack two 2D tuples to index 4D machine.
18 return m4d[*upper, *lower]
19
20IndexTaskMap tiles block
21
22# Map all regions of all tasks to GPU framebuffer.
23Region * * GPU FBMEM
24# Data layout in memory.
25Layout * * * C_order
26# Collect read-only regions used by task_4.
27GarbageCollect systolic *
28# At most 1 systolic task is active at a time.
29Backpressure systolic 1
(b) Mapper in Mapple for Cannon’s algorithm.
Figure A1. Application code in a task-based programming system and the corresponding Mapple mapper for the Cannon’s algorithm.
The mapper that faithfully implements Cannon’s algorithm organizes the processors into a 2D grid and blocks the GPUs on each node. Here, we use the decompose machine transformation twice to block the node and GPU dimensions of the machine to resolve the dimensionality mismatch between the task space and the processor space. Lines 2329 specify the mapping of data to memories. The Regent statement specifies that all data should be mapped to GPU framebuffer memory. The Layout directive declares that any region used by any task on any processor (denoted by three *s) is allocated in row-major (C-style) order, which is necessary to interface with cuBLAS. The GarbageCollect directive ensures the read-only tiles of $B$ and $C$ requested by systolic tasks are eagerly reclaimed to minimize memory usage. Finally, the Backpressure directive limits the number of active systolic tasks to 1 per processor to avoid allocating too many tiles at one time.
A.2. Baseline Greedy Algorithm
1
Input:
   $d$ # Number of processors
   $k$ # Dimensionality of the task space
2
3Function Greedy($d$, $k$):
5  factors $\leftarrow$ new int[$k$]
6  for $i\in\{0,\dots,k{-}1\}$ do
7     factors[$i$] $\leftarrow 1$
8 for $p\in$ primes do
9     $j\leftarrow$ ArgMin (factors) # Index of smallest current product
10     factors[$j$] $\leftarrow$ factors[$j$] $\times\ p$
11 Sort (factors, descending=True) # Sort for consistent ordering
12  return factors
A.3. Generalization of the Decompose Primitive
While Section 4 addresses isotropic communication across task space dimensions, the decompose primitive can be extended to anisotropic patterns, such as uneven halo widths and dimension-specific all-to-all exchanges (e.g., for data transposes). This generalization supports a much wider range of computations on block-structured grids. In both cases, only the objective function changes; the same minimization procedure from Section 4.3 still applies.
A.3.1. Anisotropic halo communication
To generalize, we extend the communication area from Section 4.2 to a volume that captures the total data exchanged between distributed machines. In isotropic cases, where halo width is uniform, volume reduces to area. For anisotropic patterns, varying halo width across dimensions must be considered. Letting $h_{n}$ denote the width in dimension $n$, the total communication volume $V$ in a $k$-dimensional space is $V=\sum_{n=1}^{k}d_{i_{n}}h_{n}\prod_{m\neq n}l_{m}$. For $l_{n}=w_{n}d_{i_{n}}$, the volume, $V$, can also be expressed as $V=\left(\sum_{n=1}^{k}\frac{h_{n}}{w_{n}}\right)\left(\prod_{m=1}^{k}l_{m}\right)$, where $\prod_{m=1}^{k}l_{m}$ is constant for a certain task space.
A.3.2. Transpose via all-to-all communication along certain dimensions
A.4. Example Mapping Functions in Mapple
Figure A2. Example functions used by the mappers for the distributed matrix-multiplication algorithms.
A.5. Case Studies on Mapper Performance
Here we present a case study highlighting the critical role of precise control over mapping strategies. Our investigation demonstrates significant performance variations resulting from minor alterations in the hierarchical_block2D function for the Cannon’s, PUMMA, and SUMMA algorithms. Specifically, we contrast the mapping function specified by the algorithm with an alternative approach integrating runtime heuristics: the runtime system dynamically assigns the index point to one of the four GPUs on the node, selecting the GPU with the least workload at runtime, rather than adhering to a predetermined distribution.
[…]
(a) Cannon’s
(b) PUMMA
(c) SUMMA
Figure A3. Throughput comparison between the original hierarchical_block2D mapping function (specified by the Cannon’s, PUMMA, and SUMMA algorithm) and a runtime heuristics-based mapper. The runtime heuristics-based mapping can lead to out-of-memory error (denoted by “OOM”).